# Learning Robotic Policy with Imagined Transition: Mitigating the Trade-off between Robustness and Optimality

### v1

@@ Abstract
Existing quadrupedal locomotion learning paradigms usually rely on extensive domain randomization to alleviate the sim2real gap and enhance robustness. It trains policies with a wide range of environment parameters and sensor noises to perform reliably under uncertainty. However, since optimal performance under ideal conditions often conflicts with the need to handle worst-case scenarios, there is a trade-off between optimality and robustness. This trade-off forces the learned policy to prioritize stability in diverse and challenging conditions over efficiency and accuracy in ideal ones, leading to overly conservative behaviors that sacrifice peak performance. In this paper, we propose a two-stage framework that mitigates this trade-off by integrating policy learning with imagined transitions. This framework enhances the conventional reinforcement learning (RL) approach by incorporating imagined transitions as demonstrative inputs. These imagined transitions are derived from an optimal policy and a dynamics model operating within an idealized setting. Our findings indicate that this approach significantly mitigates the domain randomization-induced negative impact of existing RL algorithms. It leads to accelerated training, reduced tracking errors within the distribution, and enhanced robustness outside the distribution.
@@ I INTRODUCTION
Learning-based locomotion control methods for quadruped robots have achieved remarkable results in recent years [1, 2, 3, 4, 5, 6]. To adapt to various wild environments, these methods collected large amounts of data in massive parallel simulators with various dynamics and terrains for policy training via domain randomization [7] and curriculum learning [8]. This data-driven paradigm allows RL-based robots to adapt to realistic terrain and complex behavior commands.
Domain randomization is a vital technique to ensure the policy’s robustness. However, previous research [9, 10] found that it trades optimality for robustness and leads to conservative policies. In Figure 1, we observe that locomotion policies trained using modern RL algorithms can rapidly achieve high performance in simulators with fixed dynamics, provided that robustness and the sim2real gap are not considered. When robustness is incorporated into the training process, the resulting policies tend to exhibit more conservative behaviors, leading to lower cumulative rewards during training and reduced body height during deployment, as a precaution against unforeseen perturbations. There exists a trade-off between optimality and robustness in current RL-based locomotion paradigms.
@@
This phenomenon is caused by the large amount of randomness introduced by domain randomization, which greatly increases the exploration space of policy training. In addition, this is a Partially Observable Markov Decision Process (POMDP) problem [11, 12], and the agent cannot obtain all the environmental parameters. Therefore, the behavior of policies tends to be conservative to cope with unknown disturbances. However, this conservatism leads to a decrease in training efficiency because the robot is difficult to obtain high rewards during training. Therefore, we need a method to mitigate this phenomenon to balance optimality and robustness better.
In the training paradigm of robotic policies, we always hope that the robot operates according to a certain pattern. For example, in manipulation tasks, it is desired that the robotic arm accurately follows a trajectory. In the locomotion of quadruped robots, it is desired that a specific gait (e.g., trotting, walking, running, etc.) is maintained. We hope that this behavior pattern can remain consistent when facing different environmental parameters and disturbances. For quadrupedal robots, the desired gait in undisturbed simulators is easy to obtain via RL training without dynamics randomization, but it is difficult to maintain robustness when deployed in complex, real-world environments. However, the ideal gait can be used as a reference to guide the learning of robust policies in complex environments. To achieve this goal, we propose LIT (Learning with Imagined Transition), a two-stage framework training paradigm that enables the learning of robust locomotion policies from ideal gait.
To get a gait like the one ideally trained in an undisturbed environment, we need to compress the action and observation corresponding to the optimal gait into the neural network. This stage is called reference learning in our framework. It trains an ideal policy and a dynamics model under the non-perturbative simulation environment. They provide the reference action and the imagined next observation to construct the imagined transition of the ideal gait. After that, the imagined transition should be used to guide the learning of robust policy. We utilized the imagined transition as part of the network input and optimized it via RL. This approach not only increases the training efficiency but also improves the generalization of the locomotion policy. Additionally, the predicted observation from the dynamics model may encounter Out-of-Distribution (OOD) problems due to the distribution shift between the complex and ideal environments. If we include it in the optimization objective, it will lead to training crash or non-convergence. Therefore, we designed an adjustment mechanism of the dynamics model to mitigate the negative impact of OOD observation on policy learning.
@@
$\bullet$ We propose LIT, an RL-based locomotion training paradigm that includes ideal gait learning and robust policy learning with imagined transitions.
$\bullet$ We develop a method to capture the desired gait in the simulator through joint training of the policy and dynamics model in a non-perturbative simulation environment.
$\bullet$ We introduce a dynamics model with an adjustment mechanism to mitigate the adverse effects of errors in the dynamics model when encountering OOD observations.
$\bullet$ We conduct extensive experiments and ablations in both simulators and real-world scenarios. The empirical results demonstrate that LIT improves training efficiency and enhances performance in both settings.
@@ II-A RL-based Legged Locomotion
To mitigate the uncertainty caused by partial observation, the teacher-student framework [3, 16] was proposed, enabling quadrupedal robots to adapt to unknown terrains and greatly improving locomotion flexibility and robustness. Much of the work [17, 18, 19] has continued in this structure with impressive results. Walk-These-Way [5] proposes a method to improve the generalization of quadrupedal locomotion by training the robot in diverse environments. Another attempt is introducing representational learning, such as [20, 4, 6]. DreamWaQ [4], leverages variational autoencoders (VAE) [21] to learn environmental representations, while HIMloco [6] employs contrastive learning techniques to learn terrain representations, further enhancing the robot’s adaptability to various environments. Recent advances [22, 23] introduce model-based control [24] and adaptive control [25] to improve the interpretability and robustness of RL-based locomotion. To address the challenges of the sim2real gap, domain randomization has been commonly used [1, 2, 3, 4, 6]. It is a class of methods in which the policy is trained with a wide range of environment parameters and sensor noise to learn behaviors that are robust in this range. However, domain randomization trades optimality for robustness, sometimes leading to an over-conservative policy.
@@ II-B Model-based RL
Model-based Reinforcement Learning (MBRL) distinguishes itself by incorporating a learned dynamics model of the environment, which predicts future states given the current state and action. This allows MBRL methods to plan and use collected data efficiently. Popular frameworks in this area include Dreamer series [26, 27, 28], which uses a learned world model to perform planning in latent space, and TDMPC series [29, 30], which integrates model predictive control to improve sample efficiency. In the legged locomotion tasks, recent MBRL approaches have demonstrated the power of learned dynamics models. DayDreamer [31], inspired by Dreamer, enables effective planning and control for robotic locomotion. PIP-loco [32] also adopts a model-based approach and combines it with model predictive control, resembling TDMPC. These methods utilize the learned model for trajectory planning or representation learning, leveraging the learned dynamics model to enhance performance. In contrast to these works, our approach utilizes the dynamics model not for planning or representation, but as a reference to predict the next state, thereby more directly affecting policy learning.
@@ III Preliminary
Figure 2: Framework Overview of LIT. The left part is gait reference learning in fixed dynamics. The right part is policy learning with imagined transition. $\mathbf{a}^{r}_{t}$ and $\mathbf{o}^{r}_{t+1}$ mean reference action and imagined next observation respectively. $[\mathbf{o}_{t},\mathbf{a}^{r}_{t},\mathbf{o}^{r}_{t+1}]$ constitutes the imagined transition
@@ IV-A Framework Overview
Our framework is shown in Fig. 2. It is a general approach that can be incorporated into most of the previous RL-based Locomotion methods such as RMA [3], Dreamwaq [4], HIMLoco [6], etc., and here we use HIMLoco as the backbone algorithm. It consists of two parts, Gait Reference Learning and Policy Learning with Imagined Transition. Section IV-B discusses how to perform Gait Reference Learning. It aims to obtain the optimal gait from a fixed simulation environment and compress the optimal gait into an ideal policy and a dynamics model. Section IV-C presents how to use gait reference to guide policy learning.
IV-B Gait Reference Learning in Fixed Dynamics
@@ IV-B2 Dynamics Model
To obtain reference observations, we train a dynamics model under the non-perturbative simulation environment. The output of the dynamics model includes $\mu$ and $\sigma$ in 1, which means the mean and standard deviation of the next observation estimation. We sample the observation estimation $\mathbf{\hat{o}}_{t+1}$ from normal distribution $N(\mu,\sigma)$ for network optimization. The optimization objective of the dynamics model is the maximum likelihood estimation, represented by the formula 3.
$\mu,\sigma\sim P(\cdot|\mathbf{o}_{t},\mathbf{a}_{t})$ (1)
$\mathbf{\hat{o}}_{t+1}\sim N(\mu,\sigma)$ (2)
$\mathcal{L}_{MLE}=-\mathbf{\hat{o}}_{t+1}log(\mathbf{o}_{t+1})$ (3)
@@
To prevent training and deployment collisions caused by model inaccuracies, we developed an adjustment mechanism to adjust the prediction of the next observation. According to the general law of supervised learning, the larger the standard deviation $\sigma$ the less reliable the prediction $\mu$ of the dynamics model. Thus, when the variance is large, we should minimize the effect of the observation reference on the subsequent network. Inspired by [4, 33], we multiply the weight $(1-norm(\sigma))\in[0,1]$ dot by $\mu$ with the following equation 4.
Experiments in 8 showed that this design significantly improved the robot’s performance in the unseen region.
$\mathbf{o}^{r}_{t+1}\sim f(\mu,\sigma)=(1-norm(\sigma))\cdot\mu$ (4)
@@ IV-C Policy Learning with Imagined Transition
In short, the only difference in our policy network is that we are adding the input of action and observation reference, which represents gait reference. The combination of $\mathbf{o}$, $\mathbf{a}^{r}_{t}$, and $\mathbf{o}^{r}_{t+1}$ forms a complete state transition, which is our imagined transition. As shown in the bottom of Fig.5, we concatenate reference action $\mathbf{a}^{r}_{t}$ and imagined next observation $\mathbf{o}^{r}_{t+1}$ from the gait reference with proprioceptive observations $\mathbf{o}_{t}$, whereas the baseline, depicted on the top, follows the architecture used in[4, 6].
$\mathbf{a}_{t}\sim\pi(\cdot|\mathbf{o}_{t-H:t},\mathbf{z},\mathbf{\hat{v}})$ (5)
$\mathbf{a}_{t}\sim\pi(\cdot|\mathbf{o}_{t-H:t},\mathbf{z},\mathbf{\hat{v}},\mathbf{a}^{r}_{t},\mathbf{o}^{r}_{t+1})$ (6)
Intuitively, solving in the solution space close to the prior gait information input should be more efficient than searching for the optimal solution directly in the huge solution space. This simple change has resulted in a significant increase in the overall performance of policy learning.
@@ V-A Experimental Setup
(A) Baseline: Train without observation reference and action reference, equivalent here to [6]; (B) Ours w/o Adjust: Ours without observation reference adjustment, the mean value output from the dynamics model is used directly as the observation reference; (C) Ours w Residual: Action reference is summed with the actor output as action output, similar to [34, 35]; (D) Ours w/o Action Reference: Action reference input is set to zero tensor; (E) Ours w/o State Reference: State reference input is set to zero tensor; (F) Ours;
@@ V-B1 Command Tracking
Terrain Types Velocity Types Baseline Ours
Smooth Slopes Linear 0.170 0.120
@@
Rough Slopes Linear 0.387 0.162
@@ V-B2 Learning Curve
Figure 7: Ablation studies with learning curves. From top to bottom: (1) normalized linear velocity tracking score, (2) normalized angular velocity tracking score, and (3) maximum reachable terrain level in training environments.
@@ V-B4 Adding push to robots
[…]
Figure 11: Fall resistant visualization with external push. ‘Common’ represents cases where both ours and baseline fall. ‘Ours’ and ‘Baseline’ represent cases where each falls but the other survives.
@@ V-C Evaluation on Real World
[…]
Figure 12: Comparison of command tracking performance in real-world scenarios. The results are averaged over ten trials. A lower tracking error indicates better deployed performance.
@@ V-D Implementation Details
Linear velocity tracking $\exp\left\{-\frac{\|\mathbf{v}_{xy}^{\text{cmd}}-\mathbf{v}_{xy}\|_{2}^{2}}{0.25}\right\}$ 1.0
Angular velocity tracking $\exp\left\{-\frac{\left(\omega_{\text{yaw}}^{\text{cmd}}-\omega_{\text{yaw}}\right)^{2}}{0.25}\right\}$ 0.5
Lin. velocity $(z)$ $v_{z}^{2}$2.0
Ang. velocity $(xy)$ $\|\boldsymbol{\omega}_{xy}\|_{2}^{2}$0.05
Orientation $\|\mathbf{g}\|_{2}^{2}$0.2
Joint accelerations $\|\ddot{\boldsymbol{\theta}}\|^{2}$ $-2.5\times 10^{-7}$
Joint power $|\boldsymbol{\tau}\|\dot{\boldsymbol{\theta}}|^{T}$ $-2\times 10^{-5}$
Body height $\left(h^{\text{target}}-h\right)^{2}$1.0
Foot clearance $\sum_{i=0}^{3}\left(p_{z}^{\text{target}}-p_{z}^{i}\right)^{2}\cdot v_{xy}^{i}$0.01
Action rate $\|\mathbf{a}_{t}-\mathbf{a}_{t-1}\|_{2}^{2}$0.01
Smoothness $\|\mathbf{a}_{t}-2\mathbf{a}_{t-1}+\mathbf{a}_{t-2}\|_{2}^{2}$0.01
@@
CoM $[-0.05,0.05]$ $\mathrm{m}$
Payload Mass $[-1,2]$ $\mathrm{Kg}$
Ground Friction $[0.2,1.25]$
Motor Strength [0.9,1.1]×[0.9,1.1$\times$[ 0.9 , 1.1 ] × motor torque $\mathrm{Nm}$
Joint $K_{p}$ [0.9,1.1]×[0.9,1.1$\times$[ 0.9 , 1.1 ] × 40.0 Nm/rad
Joint $K_{d}$ [0.9,1.1]×[0.9,1.1$\times$[ 0.9 , 1.1 ] × 1.0 Nms/rad
Initial Joint Positions [0.5,1.5]×[0.5,1.5$\times$[ 0.5 , 1.5 ] × nominal value $\mathrm{rad}$
@@
External Force $[-30,30]$ $\mathrm{N}$
@@
To facilitate robot learning, we adopt a terrain curriculum and command curriculum inspired by the approaches of [36], [37], and [6]. Specifically, we construct a height field map comprising 200 distinct terrains organized in a $20\times 10$ grid. Within each row, terrains are arranged in increasing order of difficulty, and each grid cell measures $10\times 10\,m^{2}$. Initially, robots are placed on the simplest terrain types. As the robot’s performance improves, the terrain difficulty level increases once the robot achieves at least 80% of the linear tracking reward. Conversely, the difficulty decreases if the robot cannot traverse half of the terrain within a single episode. For terrains involving stairs or discrete obstacles, we sample longitudinal and lateral linear velocity commands from the range $[-1.0,1.0]$ m/s and the horizontal angular velocity from $[-2.0,2.0]$ rad/s. For slopes and rough terrains, the range of longitudinal linear velocity is extended to $[-3.0,3.0]$ m/s, and the range for horizontal angular velocity is expanded to $[-3.0,3.0]$ rad/s. All commands are sampled independently for each robot from the specified ranges, with sampling occurring every 25-time steps.

### v2

@@ Abstract
Existing quadrupedal locomotion learning paradigms usually rely on extensive domain randomization to alleviate the sim2real gap and enhance robustness. It trains policies by introducing various disturbances into simulated environments to improve their adaptability under uncertainty. However, since optimal performance without disturbances often conflicts with the need to handle worst-case scenarios, this paradigm has a trade-off between optimality and robustness. This trade-off forces the learned policy to prioritize stability in diverse and challenging environments over efficiency and accuracy in no-disturbed ones, leading to overly conservative behaviors that sacrifice peak performance. Inspired by disturbance rejection control, we propose a two-stage framework that integrates policy learning with motion reference to mitigate this phenomenon. This framework enhances the conventional reinforcement learning (RL) approach by incorporating imagined transitions as reference inputs. The imagined transitions are derived from an optimal policy and a dynamics model operating within an idealized setting without disturbances. By providing this reference, we found that the policy could generate actions that are better aligned with the desired motion, which led to accelerated training and reduced tracking errors under external disturbances.
@@ I INTRODUCTION
Learning-based locomotion control methods for quadruped robots have achieved remarkable results in recent years [1, 2, 3, 4, 5, 6]. To adapt to various wild environments, these methods collected large amounts of data in massive parallel simulators with extensive dynamics variations via domain randomization [7]. This data-driven paradigm allows RL-based robots to adapt to realistic terrain and unseen disturbances.
While domain randomization is a widely used technique to improve the robustness of robot policies, previous studies [8, 9] have shown that it often sacrifices optimality, resulting in conservative behaviors. For instance, when trained with large payload variations, the learned policy tends to maintain a lower base height rather than achieve the target height specified in the reward function. This example reveals that the robustness achieved through domain randomization essentially comes at the cost of compromising certain task objectives under unknown changes. Consequently, it becomes difficult to consistently meet high task standards, which limits the applicability of domain randomization in complex tasks, particularly those requiring high accuracy.
@@
To mitigate this problem, we require a method that can maintain task standards without compromising adaptability under disturbances. A key element for enabling robots to recover from unknown perturbations is the explicit identification of the goal that the policy should adapt to. This aligns with the control-theoretic perspective, particularly classical disturbance rejection control, where reference signals are used to compute counteractions that cancel perturbation effects These reference signals define the desired system behavior and represent the performance the controller seeks to preserve despite disturbances. In RL policy training, however, such references are only implicitly encoded in the reward functions, and no explicit goal-oriented guidance is provided. Inspired by disturbance rejection control, we therefore propose to supply explicit motion references in RL policy learning. These references not only provide a clearer objective for training, enabling the policy to learn to resist disturbances and track desired motions, but also deliver richer guidance for policy inference.
Normally, it is straightforward to train a policy in simulation under a fixed set of environmental parameters without any randomization. Although such a policy cannot be directly deployed in reality due to limited robustness, it can still serve as a reference motion to guide both policy training and inference. Alternatively speaking, a well-trained policy in an ideal simulation can be leveraged to generate goal information that assists the deployed policy in making better decisions. To this end, we propose LIT (Learning with Imagined Transition), a two-stage framework training paradigm that enables the learning of robust locomotion policies from ideal motion. By introducing “ideal” actions and transitions, the policy could generate actions that are better aligned with nominal dynamics under disturbances (as illustrated by the gray line in Fig. 1), thereby enabling more robust and accurate task execution.
Specifically, we reconstruct a learnable dynamics model to approximate transition propagation, obtained from a non-perturbative simulation environment together with an ideal policy. These provide reference actions and next-step ideal observations for policy training and inference, guiding the policy toward robust locomotion even under external disturbances. To achieve this, we incorporate ideal transitions into the policy input while continuing to optimize the policy in an RL manner. This approach not only improves training efficiency but also enhances the generalization capability of the locomotion policy. However, the predicted observation from the dynamics model and ideal policy may become ineffective in Out-of-Distribution (OOD) scenarios. If we include it in the optimization objective, it may lead to a training crash or non-convergence. To address this, we designed an adjustment mechanism of the dynamics model to mitigate the negative impact of OOD observation during policy learning.
@@
$\bullet$ We propose LIT, a two-stage RL-based locomotion paradigm that consists of motion reference learning and robust policy learning, leveraging ideal transitions obtained through joint training of the policy and dynamics model in a non-perturbative simulation environment.
$\bullet$ We design an adjustment mechanism to mitigate the adverse effects of dynamics model errors when encountering out-of-distribution (OOD) observations, ensuring efficient and stable policy training.
$\bullet$ We conduct extensive experiments and ablations in both simulators and real-world scenarios. The empirical results demonstrate that LIT improves training efficiency and significantly enhances quadruped locomotion performance under unknown disturbances.
@@ II-A RL-based Legged Locomotion
To mitigate the uncertainty caused by partial observation, the teacher-student framework [3, 13] was proposed, enabling quadrupedal robots to adapt to unknown terrains and greatly improving locomotion flexibility and robustness. Much of the work [14, 15, 16] has continued in this structure with impressive results. Walk-These-Way [5] proposes a method to improve the generalization of quadrupedal locomotion by training the robot in diverse environments. Another attempt is introducing environmental representation, such as [17, 4, 6]. DreamWaQ [4], leverages variational autoencoders (VAE) [18] to learn environmental representations, while HIMloco [6] employs contrastive learning techniques to learn terrain representations. Recent advances [19, 20] introduce model-based control [21] and adaptive control [22] to improve the interpretability and robustness of RL-based locomotion. To address the challenges of the sim2real gap, domain randomization has been commonly used [1, 2, 3, 4, 6]. It is a class of methods in which the policy is trained with a wide range of environment parameters and sensor noise to learn behaviors that are robust in this range. However, domain randomization trades optimality for robustness, sometimes leading to an over-conservative policy.
@@ II-B Model-based RL
Model-based Reinforcement Learning (MBRL) incorporates a learned dynamics model of the environment, which predicts future states given the current state and action. This allows MBRL methods to plan and use collected data efficiently. Popular frameworks include Dreamer series [23, 24, 25], which uses a learned world model to perform planning in latent space, and TDMPC series [26, 27], which integrates model predictive control to improve sample efficiency. In the legged locomotion tasks, recent MBRL approaches have demonstrated the power of learned dynamics models. DayDreamer [28], inspired by Dreamer, enables effective planning and control for robotic locomotion. PIP-loco [29] also adopts a model-based approach and combines it with model predictive control, resembling TDMPC. These methods utilize the learned model for trajectory planning or representation learning, leveraging the learned dynamics model to enhance performance. In contrast to these works, our approach utilizes the dynamics model not for planning or representation, but as a reference to predict the next state, thereby more directly affecting policy learning.
@@ III Preliminary
Figure 2: Framework Overview of LIT. The left part is motion reference learning in fixed dynamics. The right part is policy learning with imagined transition. $\mathbf{a}^{r}_{t}$ and $\mathbf{o}^{r}_{t+1}$ mean reference action and imagined next observation respectively. $\mathbf{o}_{t},\mathbf{a}^{r}_{t},\mathbf{o}^{r}_{t+1}]$ constitutes the imagined transition
@@ IV-A Framework Overview
Our framework is shown in Fig. 2. It is a general approach that can be incorporated into most of the previous RL-based Locomotion methods such as RMA [3], Dreamwaq [4], HIMLoco [6], etc., and here we use HIMLoco as the backbone algorithm. It consists of two parts: Motion Reference Learning and Policy Learning with Imagined Transition. Section IV-B discusses how to perform Motion Reference Learning. It aims to obtain the optimal gait from a fixed simulation environment and compress the optimal gait into an ideal policy and a dynamics model. Section IV-C presents how to use motion reference to guide policy learning.
IV-B Motion Reference Learning in Fixed Dynamics
@@ IV-B2 Dynamics Model
To obtain reference observations, we train a dynamics model under the non-perturbative simulation environment. The output of the dynamics model includes $\mu$ and $\sigma$ in eq. 1, which means the mean and standard deviation of the next observation estimation. We use the dynamics model to output $\mu$ and $\sigma$, which parameterize a normal distribution. The training objective is to maximize the likelihood of the true next observation $\mathbf{o}_{t+1}$ under this distribution.
$\mu,\sigma=d_{\theta}(\cdot|\mathbf{o}_{t},\mathbf{a}_{t})$ (1)
$\mathcal{L}_{MLE}=-\mathbb{E}[\log p(\mathbf{o}_{t+1}|\mu,\sigma)]$ (2)
@@
To prevent training and deployment collisions caused by model inaccuracies, we developed an adjustment mechanism to adjust the prediction of the next observation in the following robust policy learning stage (see Sec. IV-C) and deployment. According to the general law of supervised learning, the larger the standard deviation $\sigma$ the less reliable the prediction $\mu$ of the dynamics model. Thus, when the variance is large, we should minimize the effect of the observation reference on the subsequent network. Inspired by [4, 30], we multiply the weight $(1-norm(\sigma))\in[0,1]$ dot by $\mu$ with the following equation 3. Experiments in 8 showed that this design significantly improved the robot’s performance in the unseen region.
$\mathbf{o}^{r}_{t+1}=(1-norm(\sigma))\cdot\mu$ (3)
@@ IV-C Policy Learning with Imagined Transition
In short, the only difference in our policy network is that we are adding the input of action and observation reference, which represents motion reference. The combination of $\mathbf{o}$, $\mathbf{a}^{r}_{t}$, and $\mathbf{o}^{r}_{t+1}$ forms a complete state transition, which is our imagined transition. As shown in the bottom of Fig.5, we concatenate reference action $\mathbf{a}^{r}_{t}$ and imagined next observation $\mathbf{o}^{r}_{t+1}$ from the motion reference with proprioceptive observations $\mathbf{o}_{t}$, whereas the baseline, depicted on the top, follows the encoder-policy network architecture used in[4, 6].
$\mathbf{a}_{t}\sim\pi(\cdot|\mathbf{o}_{t-H:t})$ (4)
$\mathbf{a}_{t}\sim\pi(\cdot|\mathbf{o}_{t-H:t},\mathbf{a}^{r}_{t},\mathbf{o}^{r}_{t+1})$ (5)
Intuitively, solving in the solution space close to the motion prior should be more efficient than searching for the optimal solution directly in the huge solution space. This simple change has resulted in a significant increase in the overall performance of policy learning.
@@ V-A Experimental Setup
(A) Baseline: Train with current and history observations, equivalent here to [6]; (B) Ours w/o Adjust: Ours without observation reference adjustment, the mean value output from the dynamics model is used directly as the observation reference; (C) Ours w Residual: Action reference is summed with the actor output as action output, similar to residual policy in [31, 32]; (D) Ours w/o Action Reference: Action reference input is set to zero tensor; (E) Ours w/o State Reference: State reference input is set to zero tensor; (F) Ours;
@@ V-B1 Command Tracking
Terrain Types
Velocity Types
Baseline Ours
Smooth Slopes
Linear 0.170 0.120
@@
Rough Slopes
Linear 0.387 0.162
@@ V-B2 Learning Curve
Figure 7: Ablation studies with learning curves. From top to bottom: (1) normalized linear velocity tracking score, (2) normalized angular velocity tracking score, and (3) maximum reachable terrain level in training.
@@ V-B4 Adding push to robots
Pushing a robot with external force is a widely used method for evaluating the robustness of its locomotion. In this experiment, the push force was applied by adding a random velocity to the quadruped’s base along the $xy$-plane of the world frame. The robot was instructed to walk forward with a command of $1,0,0]\,\text{m/s}$, while simultaneously responding to the unpredicted random push. The magnitude of the added velocity determines the intensity of the push, with larger magnitudes corresponding to more forceful pushes. To assess the robustness of our approach in comparison to the baseline, pushes were randomly sampled 2000 times within a velocity range of $3,3]\,\text{m/s}$. The results in Figure 10 illustrate that our method outperforms the baseline in terms of fall resistance when subjected to external pushes.
Figure 11 provides a detailed visualization of the standing versus falling status under various push magnitudes and directions for different algorithms. Given the command $1,0,0]\,\text{m/s}$, it can be observed that Figure 11 exhibits near-symmetry along the $x$-axis, but not along the $y$-axis. All methods include push perturbations within the range $1,1]\,\text{m/s}$ during the robust training phase, and consequently, they perform well within this range. However, as the perturbation range increases, the advantage of our approach becomes increasingly pronounced. Our method outperforms the baseline by surviving more than 300 additional times in 2000 randomized tests.
Figure 11: Fall resistant visualization with external push. ‘Both’ represents cases where both ours and baseline fall. ‘Ours’ and ‘Baseline’ represent cases where each falls but the other survives.
@@ V-C Evaluation on Real World
To test the effectiveness of policy deployment in the real world, we designed a variety of challenging scenarios as shown in Figure 9. We deployed the policy for training 2000 iterations on the UniTree A1 robot, with the PD controller’s parameters set to $Kp=40.0$ and $Kd=1.0$. The robot was instructed to walk forward with a command of $1,0,0]\,\text{m/s}$ in five scenarios. We measured the tracking errors $v_{x,y}-v^{target}_{x,y}||^{2}$ as the performance metric to evaluate the robot’s ability to track the linear velocity command, where the body velocity $v_{x,y}$ is collected by the robot’s IMU. The results in Figure 12 show that our method has a lower tracking error in five scenarios compared to the baseline.
Figure 12: Comparison of command tracking performance in real-world scenarios. The results are averaged over ten trials. Error bars represent standard error. A lower tracking error indicates better deployed performance.
@@ V-D Implementation Details
Linear velocity
tracking
$\exp\left\{-\frac{\|\mathbf{v}_{xy}^{\text{cmd}}-\mathbf{v}_{xy}\|_{2}^{2}}{0.25}\right\}$ 1.0
Angular velocity
tracking
$\exp\left\{-\frac{\left(\omega_{\text{yaw}}^{\text{cmd}}-\omega_{\text{yaw}}\right)^{2}}{0.25}\right\}$ 0.5
Lin. velocity $(z)$ $v_{z}^{2}$ -2.0
Ang. velocity $(xy)$ $\|\boldsymbol{\omega}_{xy}\|_{2}^{2}$ -0.05
Orientation $\|\mathbf{g}\|_{2}^{2}$ -0.2
Joint accelerations $\|\ddot{\boldsymbol{\theta}}\|^{2}$ $2.5\times 10^{-7}$
Joint power $|\boldsymbol{\tau}\|\dot{\boldsymbol{\theta}}|^{T}$ $2\times 10^{-5}$
Body height $\left(h^{\text{target}}-h\right)^{2}$ -1.0
Foot clearance $\sum_{i=0}^{3}\left(p_{z}^{\text{target}}-p_{z}^{i}\right)^{2}\cdot v_{xy}^{i}$ -0.01
Action rate $\|\mathbf{a}_{t}-\mathbf{a}_{t-1}\|_{2}^{2}$ -0.01
Smoothness $\|\mathbf{a}_{t}-2\mathbf{a}_{t-1}+\mathbf{a}_{t-2}\|_{2}^{2}$ -0.01
@@
CoM [0.05,0.05][-0.05,0.05] $\mathrm{m}$
Payload Mass [1,2][-1,2] $\mathrm{Kg}$
Ground Friction $[0.2,1.25]$ -
Motor Strength $0.9,1.1]\times$ motor torque $\mathrm{Nm}$
Joint $K_{p}$ $0.9,1.1]\times$ 40.0 Nm/rad
Joint $K_{d}$ $0.9,1.1]\times$ 1.0 Nms/rad
Initial Joint Positions $0.5,1.5]\times$ nominal value $\mathrm{rad}$
@@
External Force [30,30][-30,30] $\mathrm{N}$
@@
To facilitate robot learning, we adopt a terrain curriculum and command curriculum inspired by the approaches of [33], [34], and [6]. Specifically, we construct a height field map comprising 200 distinct terrains organized in a $20\times 10$ grid. Within each row, terrains are arranged in increasing order of difficulty, and each grid cell measures $10\times 10\,m^{2}$. Initially, robots are placed on the simplest terrain types. As the robot’s performance improves, the terrain difficulty level increases once the robot achieves at least 80% of the linear tracking reward. Conversely, the difficulty decreases if the robot cannot traverse half of the terrain within a single episode. For terrains involving stairs or discrete obstacles, we sample longitudinal and lateral linear velocity commands from the range [1.0,1.0][-1.0,1.0] m/s and the horizontal angular velocity from [2.0,2.0][-2.0,2.0] rad/s. For slopes and rough terrains, the range of longitudinal linear velocity is extended to [3.0,3.0][-3.0,3.0] m/s, and the range for horizontal angular velocity is expanded to [3.0,3.0][-3.0,3.0] rad/s. All commands are sampled independently for each robot from the specified ranges, with sampling occurring every 25-time steps.